# Identifying Sub-networks in Neural Networks via Functionally Similar Representations

### v1

@@ Abstract
Mechanistic interpretability aims to provide human-understandable insights into the inner workings of neural network models by examining their internals. Existing approaches typically require significant manual effort and prior knowledge, with strategies tailored to specific tasks. In this work, we take a step toward automating the understanding of the network by investigating the existence of distinct sub-networks. Specifically, we explore a novel automated and task-agnostic approach based on the notion of functionally similar representations within neural networks, reducing the need for human intervention. Our method identifies similar and dissimilar layers in the network, revealing potential sub-components. We achieve this by proposing, for the first time to our knowledge, the use of Gromov-Wasserstein distance, which overcomes challenges posed by varying distributions and dimensionalities across intermediate representations—issues that complicate direct layer-to-layer comparisons. Through experiments on algebraic and language tasks, we observe the emergence of sub-groups within neural network layers corresponding to functional abstractions. Additionally, we find that different training strategies influence the positioning of these sub-groups. Our approach offers meaningful insights into the behavior of neural networks with minimal human and computational cost.
@@ 1 Introduction
Rapid progress in transformer language models has directed attention towards understanding the underlying causes of new capabilities. Like many other neural methods, large language models (LLMs) are mostly black-box models and explainable artificial intelligence aims to offer insights and improve human understanding of these LLMs. Recently, mechanistic interpretability research has gained popularity, focusing on reverse-engineering models into human-understandable algorithms, using methods such as computational graphs and circuits [30, 5]. Examples include recovering internal mechanisms of LLMs to solve typical mathematical problems such as modular sum [45].
@@
Current approaches primarily rely on extensive manual inspection and trial-and-error to reverse engineer networks by discovering a sequence of learned functions that produce a desired output. This process requires significant human prior knowledge with strategies tailored to specific tasks such as algebraic problems [4]. Automatic discovery of these functions is a difficult task due to the large search space involved. We investigate a simpler task that can be considered as an initial step towards automating the discovery of these functions: is it possible to detect how many distinct (complex) functions exist in a learned network, and which layers correspond to each such function? In this paper, we take such a step towards a more automatic and task-agnostic approach to mechanistic interpretability by identifying sub-components in neural networks that are functionally different from each other. Specifically, we treat the intermediate layers of neural networks as computing different functions of the input and base our approach on functional similarity. These distinctive subnetworks would represent one function in a sequence of different ones, and knowing such a partition would facilitate further decoding these functions leading a better understanding of the network.
The nature of functional similarities differ depending on which of the two cases is true: either we have a hypothesized target function that part of the network may (approximately) compute, or we lack such a hypothesis. In the case of a given target function, functional similarity relates to the output of the target function, and the problem is one of search to find a network layer that best approximates it. This is akin to comparing two functions based on the values they take for the same input samples. The use of probes attached to representations is a popular method for detecting such similarities.
@@
Figure 1: Overview of our approach where we use representations to identify functionally distinct sub-networks (darker blocks) leveraging GW distance.
@@
The more realistic case with neural networks is the absence of a target function, which is the primary focus of this paper. In this case, we propose to measure the functional similarity between layers of the network, and the problem is to identify distinctive sub-components within the network by finding which layers are functionally less similar to previous layers as shown in Figure 1. The idea is that less similar layers (brighter colors) may indicate boundaries of a sub-component, while more similar layers (darker colors) likely belong to the same sub-component. While it may be tempting to also use probes here for layer comparisons, the lack of targets makes it unclear how the representations in these layers should be transformed by the probes for comparison. We need a more direct way to compare these representations. Hence, we propose to measure similarity using Gromov-Wasserstein (GW) distance [44] between representations from different layers of the network. As elaborated further in Section 4.1, GW allows distance computation between distributions supported on two different metric spaces with different supports and potentially different dimensions, which is common across different layers in neural networks. GW is also invariant to permutation of the representation within a layer, a crucial property since neural networks are known to have permutation symmetries [15]. As such, GW can effectively identify genuinely distinct behaviors across (groups of) layers.
@@
We validate our approach on algebraic and NLP sentiment analysis tasks, showing that GW distance provides a systematic way to analyze and identify subnetworks. Additionally, our findings provides a holistic view on differences in representations of models trained with different strategies. We observe clear patterns in the form of block structures among different layers, suggesting there exist sub-networks that have different functions, particularly at the transition layers where major functional changes may occur. Moreover, the GW distance can also be used to observe the emergence of subnetworks during training process. Overall, we hope that our method can improve the efficiency of mechanistic interpretation by finding subnetworks in larger models, reducing the need for extensive human effort and potentially contributing to a further understanding of neural network behaviors.
@@ 2 Background and Related Work
With its popularity, mechanistic interpretability has become a disparate area and we survey various directions within it.
Algorithm discovery in algebraic problems Modular sum algorithms [30] are studied in the context of progress measure in emerging abilities of transformers. The authors in [45] show, via various inspection such as gradient symmetry, logit patterns, and input pair relations, that one class of algorithm is preferred than another. Other math problems such as the greatest common divisor [4] also show insights into the inner mechanisms of neural networks.
Other types of discovery Automated circuit discovery [5, 36] intends to find a computational graph that is far sparser without sacrificing performance, using KL measures. Heuristic Core [1] hypothesizes that there exists a set of attention heads that encompasses all subnetworks, learning shallow and non-generalizable features. Generalization is achieved by incorporating additional attention heads with higher-level features. Average gradient outer product [35] uses eigen directions of weight matrices to identify task-relevant gradient directions. Moreover, various work have explored neuron semantics and possible disentanglement of semantics [2, 11, 18], as well as concept representations [33].
Real applications An intervention study [32] comparing clean and corrupted images via causal tracing shows that vision and language components are processed independently in the intermediate layers, interacting only in the final layers. [31] studies factual and counterfactual statements, investigating how they compete with each other through their differences in attention logits and weights. [16] finds that relation decoding can be represented by a linear relation embedding with causal semantics. [43] studies the hallucination problem in factual completion tasks.
@@
Methods of Studying Mechanisms Weight inspection and manipulation are commonly employed techniques to gain insights into the inner work of networks, including studying periodicity [30] and weight gradients [45]. Modifying or ablating the activation of a specific model components [19, 22], including attention knockout [42], and even direct modification of attention matrix [31, 14] are prevalent. Another popular approach involves representation and output inspection [28], including logit patterns [45], residual stream [31], and periodicity [30]. Causal mediation analysis is used to compute the indirect treatment effect [28, 43] with perturbed embedding, often with Gaussian noises. Rather than just inspection of outputs, many works have proposed to map the output to some target and is a popular technique for analyzing how neural activations correlate with high-level concepts [18]. Linear probes are generally used. [17] uses a nonparametric probe ($k-$nearest-neighbor) to classify outputs for reasoning tasks.
@@ 3 Functional Similarity Within Neural Networks
We aim to identify sub-components of a neural network based on their mechanistic functions. When we have a hypothesized target function that the sub-component may compute, this can be formulated as a similarity search problem In this context, we search for candidate representations within neural networks’ outputs that represents changes in function computation. The search problem consists of three key elements: the search space, the search target, and the similarity measures used to evaluate how closely the candidates in the search space match the target.
@@ 3.1 Similarity Measures
Let $f_{0}$ be a function that produces a set of output $Y^{0}$ given a set of input $X$, where $Y^{0}\in{\mathbb{R}}^{n\times d_{y}}$ and $X\in{\mathbb{R}}^{n\times d_{x}}$ are assumed to have dimensions $d_{y}$ and $d_{x}$, respectively, with $n$ being the set size, without loss of generality. Let $f$ be another function that produces $Y$ given $X$.
Functional Similarity. Similarity between two functions, $f_{0}$ and $f$, over a set of input $X$, can be measured by the similarity between their outputs $Y_{0}=f_{0}(X)$ and $Y=f(X)$.

### v2

@@ Abstract
Providing human-understandable insights into the inner workings of neural networks is an important step toward achieving more explainable and trustworthy AI. Existing approaches to such mechanistic interpretability typically require substantial prior knowledge and manual effort, with strategies tailored to specific tasks. In this work, we take a step toward automating the understanding of the network by investigating the existence of distinct sub-networks. Specifically, we explore a novel automated and task-agnostic approach based on the notion of functionally similar representations within neural networks to identify similar and dissimilar layers, revealing potential sub-networks. We achieve this by proposing, for the first time to our knowledge, the use of Gromov-Wasserstein distance, which overcomes challenges posed by varying distributions and dimensionalities across intermediate representations—issues that complicate direct layer-to-layer comparisons. On algebraic, language, and vision tasks, we observe the emergence of sub-groups within neural network layers corresponding to functional abstractions. Through downstream applications of model compression and fine-tuning, we show the proposed approach offers meaningful insights into the behavior of neural networks with minimal human and computational cost.
@@ 1 Introduction
Rapid progress in transformer language models has directed attention towards understanding the underlying causes of new capabilities. Like many other neural methods, large language models (LLMs) are mostly black-box models and explainable artificial intelligence aims to offer insights and improve human understanding of these LLMs. Recently, mechanistic interpretability research has gained popularity, focusing on reverse-engineering models into human-understandable algorithms, using methods such as computational graphs and circuits [6].
@@
Current approaches primarily rely on extensive manual inspection and trial-and-error to reverse engineer networks by discovering a sequence of learned functions that produce a desired output. This process requires significant human prior knowledge with strategies tailored to specific tasks such as algebraic problems [5]. Automatic discovery of these functions is a difficult task due to the large search space involved. We investigate a simpler task that can be considered as an initial step towards automating the discovery of these functions: is it possible to detect how many distinct (complex) functions exist in a learned network, and which layers correspond to each such function? Understanding neural networks through the identification of subnetworks is essential due to the complexity and opacity of modern deep learning models. Neural networks, especially those with many layers and parameters, often exhibit behaviors that are difficult to interpret holistically. Identifying subnetworks allows us to break down the model into smaller, more interpretable units, providing insights into how individual components contribute to the model’s overall performance.
To address the above question, we propose an automatic and task-agnostic approach to identify sub-components in neural networks that are functionally different from each other. Specifically, we treat the intermediate layers of neural networks as computing different functions of the input and base our approach on comparing functional similarity. These distinctive subnetworks would represent one function in a sequence of different ones, and knowing such a partition could facilitate further decoding these functions leading a better understanding of the network.
@@
Figure 1: Overview of our approach, where we use representations from different neural network layers to identify functionally distinct sub-networks (darker blocks) leveraging GW distance. The figure is an illustration based on a single transformer block, and the proposed technique can be applied to other types of network structures.
The nature of functional similarities differ depending on which of the two cases is true: either we have a hypothesized target function that part of the network may (approximately) compute, or we lack such a hypothesis. In the case of a known target function, functional similarity relates to the output of a target function, and the problem is one of search to find a network layer that best approximates it. The use of probes attached to representations is a popular method for detecting such similarities.
@@
The more realistic case with neural networks is the absence of a target function, which is the primary focus of this paper. In this case, we propose to measure the functional similarity between layers of the network, which is akin to comparing two functions based on the values they take for the same input samples. The problem then becomes to identify distinctive sub-networks by finding which layers are functionally less similar to previous layers as shown in Figure 1. The idea is that less similar layers (brighter colors) may indicate boundaries of a sub-component, while more similar layers (darker colors) likely belong to the same sub-component. While it may be tempting to also use probes here for layer comparisons, the lack of targets makes it unclear how the representations in these layers should be transformed by the probes for comparison. We need a more direct way to compare these representations.
We propose to measure function similarity using Gromov-Wasserstein (GW) distance [53] between representations from different layers of the network. As elaborated further in Section 4.1, GW allows distance computation between distributions supported on two different metric spaces with different supports and potentially different dimensions, which is common across different layers in neural networks. GW is also invariant to permutation of the representation within a layer, a crucial property since neural networks are known to have permutation symmetries [16]. As such, GW can effectively identify genuinely distinct behaviors across (groups of) layers.
@@
We validate our approach on algebraic, language, and vision tasks, showing that GW distance provides a systematic way to analyze and identify subnetworks. Additionally, our findings provides a holistic view on differences in representations of models trained with different strategies. We observe clear patterns in the form of block structures among different layers, suggesting there exist sub-networks that have different functions, particularly at the transition layers where major functional changes may occur. Moreover, our approach is applicable to various downstream tasks, such as tracking the emergence of subnetworks during training process (§ 5.3) and identifying potentially redundant layers in model compression and fine-tuning (§ 5.2) Overall, our method can improve the efficiency of mechanistic interpretation by finding subnetworks in larger models, reducing the need for extensive human effort and contributing to a further understanding of neural network behaviors.
@@ 2 Background and Related Work
With its popularity, mechanistic interpretability has become a disparate area, with many different applications in vision [38], and language [37, 19, 52], and we survey a few closely related directions within it.
Algorithm discovery in algebraic problems Modular sum algorithms [35, 54] and related math problems [18, 5] are often studied in mechanistic interpretability due to their precise function definitions. These works have shown that there may exist many different sub-functions in the computation. However, most of these works require extensive human effort to reverse engineering these algorithms. We investigate whether we can automate the discovery process by studying a simpler problem of sub-function detection.
Subnetwork and Neuron Semantics Automated circuit discovery [6, 42] aims to find a computational graph that is much more sparse without sacrificing performance. Heuristic core [1] hypothesizes that there exists a set of attention heads that encompasses all subnetworks, learning shallow and non-generalizable features. Moreover, Various work have explored neuron semantics and possible disentanglement of semantics  [3, 12, 21] and concepts [39]. Instead of studying circuits and neurons, we investigate differences among neural network layers as a whole, based on an existing line of work [35]. Related, block structures within neural network layers have been observed in previous studies [36].
@@
Methods of Studying Mechanisms Weight inspection and manipulation [37, 33] are commonly employed techniques to gain insights into the inner work of networks, including studying periodicity [35] and weight gradients [54]. Modifying or ablating the activation of a specific model components [22, 25], including attention knockout [51], and even direct modification of attention matrix [37, 15] are prevalent. Another popular approach involves representation and output inspection [33], including logit patterns [54], residual stream [37], and periodicity [35]. Causal mediation analysis is used to compute the indirect treatment effect [33, 52] with perturbed embedding. Rather than just inspection of outputs, many works have proposed to map the output to some target and is a popular technique for analyzing how neural activations correlate with high-level concepts [21, 20]. Linear probes are generally used. [20] uses a nonparametric probe ($k-$nearest-neighbor) to classify outputs for reasoning tasks. Here we instead focus on studying similarity measure between layers.
Similarity Measure between Neural Network Layers There are studies that quantify the similarity between different groups of neurons [23], typically layers [11, 52], to compare different neural networks. Generally a normalized representation with desired properties is used to compare different transformer blocks, such as invariance to invertible linear transformation in canonical correlation analysis [34], orthogonal transformation, isotropic scaling, and different initializations in centered kernel alignment [24]. Other measures include representational similarity analysis [31], which studies all pairwise distances across different inputs. Wasserstein distance has been explored in measuring similarities in the context of neural networks [13, 4, 29], but they assume that different layer representations belong to the same metric space, which is very unlikely even if they have the same dimensionality as the semantics captured by each layer are likely to differ significantly. Several similarity measures [46, 7] are related to, but not exactly equivalent to, GW distance. While GW distance has been used for model merging as a regularization [43, 44], it has not been fully explored in the area of mechanistic interpretability, particularly for the subnetworks identification.
@@ 3 Functional Similarity Within Neural Networks
We aim to identify sub-components of a neural network based on their mechanistic functions. When we have a hypothesized target function that the sub-component may compute, this can be formulated as a similarity search problem. The search problem consists of three key elements: the search space, the search target, and the similarity measures used to evaluate how closely the candidates in the search space match the target. We first discuss the notation of function similarity as the measure for searching and the search space for candidates. We then discuss specific choices on the similarity measure, given whether the search target is known or not.
@@ 3.1 Similarity Measures
Let $f:X\rightarrow Y$ be a function that map $x$ in a set of input $X=\{x_{i}:x_{i}\in{\mathbb{R}}^{d_{x}}\}_{i=1}^{n}$ to $y$ in a set of output $Y=\{y_{i}:y_{i}\in{\mathbb{R}}^{d_{y}}\}_{i=1}^{n}$. Each element in $Y$ and $X$ are assumed to be a vector with dimensions $d_{y}$ and $d_{x}$, respectively, with $n$ being the set size, without loss of generality. Let $f_{0}:X\rightarrow Y^{0}$ be another function that produces $Y^{0}=\{y^{0}_{i}:y^{0}_{i}\in{\mathbb{R}}^{d_{y}}\}_{i=1}^{n}$ given $X$. Note sets can be concatenated into matrix forms as $Y^{0},Y\in{\mathbb{R}}^{n\times d_{y}}$ and $X\in{\mathbb{R}}^{n\times d_{x}}$.
Functional Similarity. Similarity between two functions, $f_{0}$ and $f$, over a set of input $X$, can be measured by the similarity between their output sets $Y_{0}=f_{0}(X)$ and $Y=f(X)$.